A box of 1L capacity is divided into two equal compartments by a thin partition which are filled with 2g H2 and 16g CH4 respectively. The pressure in each compartment is recorded as P atm. The total pressure when partition is removed will be-
A) P
B) 2P
C) P/2
D) P4

Grade:12

1 Answers

Arun
5 years ago
n(H2)=1 and n(CH4) = 1 

P(H2) = (1/(1+1))*P =P/2 
P(H2) = P/2 

P(CH4) = 1/(1+1)P = P/2 

so 
p(tot) = P 
 
Regards
Arun (askIITians forum expert)